OpenStudy (anonymous):

Can you describe and fix what I did wrong here? ax+b=0 ax=b x= b/a and the same thing goes for this one: c= ax-bx c= (a-b)x c(a-b)=x

OpenStudy (anonymous):

or better yet, \[3x+2=0\]?

OpenStudy (anonymous):

Thanks for helping but I'm still confused.

OpenStudy (anonymous):

ok lets solve \[3x+2=0\] step one: subtract \(2\) and get \[3x=-2\] step 2) divide by 3 and get \[x=-\frac{2}{3}\]

OpenStudy (anonymous):

do the same thing here \[ax+b=0\] step one: subtract \(b\) and get \[ax=-b\] (you missed the minus sign) step 2) divide by \(a\) \[x=-\frac{b}{a}\]

OpenStudy (anonymous):

clear or no? if not maybe i can explain more if so we can go on to the next one, because you were close

OpenStudy (anonymous):

Ohh thanks:) Could you help me with the second one? Oh and I woildn't have got that without you! Haha;) Thanks a lot.

OpenStudy (anonymous):

ok your first two steps were good \[c= ax-bx \\ c= (a-b)x\]

OpenStudy (anonymous):

but to get \(x\) by itself, you do not multiply by \(a-b\) but rather you need to DIVIDE by \(a-b\)

OpenStudy (anonymous):

Okay.. Yay:)

OpenStudy (anonymous):

Ohhhh

OpenStudy (anonymous):

just like if you had \[4=10x\] you would get \(\frac{4}{10}=x\)

OpenStudy (anonymous):

Thanks a bunches!! Hey how old are you anyway? You're really smart!

OpenStudy (anonymous):

when i say "divide" i don't really mean "do a division" but just write it \[c=(a-b)x\\ \frac{c}{a-b}=x\]

OpenStudy (anonymous):

Yep I got that thanks!

OpenStudy (anonymous):

ok good you familiar with black pepper?

OpenStudy (anonymous):

Yes, why?

OpenStudy (anonymous):

i am older than black pepper

OpenStudy (anonymous):

BAHAHAH I THOUGHT YOU MEANT THE SALT AND PEPPER WAY!

OpenStudy (anonymous):

lol